• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name W. Deconinck

    Log entry time 11:41:39 on October 24, 2009

    Entry number 296446

    keyword=Status of the magnets (info from Jack)

    Jack just called in, and seemed pretty optimistic about the Q2L power
    supply. He thinks it might 'just' be a problem with the lead flows which
    could maybe be fixed by changing the valve settings manually (while
    ramping up). In the Q2L a sense lead is broken, and Jack had to modify
    the board inside to PS to take that into account. It might be a bit more
    sensitive to lead flows because of this.

    Heidi and Ed are going to go down into the hall this morning to work on
    the lead flows, level meters, and magnets.